Louisiana State University Health Sciences Center Administration & Finance
 

Wednesday, January 16, 2008

1542 Tulane Ave.:

  • Siemens finished trouble-shooting the fire alarm wiring in the basement, 8th Floor, and the mechanical and animal care spaces on the roof.
    • The contractor is pulling the main fire alarm communication trunk/riser through the pipe chase from the 1st Floor to the 8th Floor.  
  •  Stratos Elevator continues to install the buffers for Elevator #8.
    • They are also installing new door hangers for Elevator #1. 
  • ATT/Bell continues to work on relocating the telephone demark to the 2nd Floor in Room 243. 
  •  Zimmer-Eschette is encapsulating the ducts on the 5th Floor. 
  • Zimmer completed the removal of floor tile on the 6th Floor and has started working on the tile removal on the 4th Floor. 
  • The painting contractor continues to float the walls in Suite 123 and Room 866.
    • The contractor continues to paint the walls on the 8th floor. 
  • Trison Constructors finished installing sheetrock for the new network closet in Room 356, 357, and 357A.  
  • Sample Results:
    • The 6th floor passed TEM clearance for asbestos floor tile removal. 
    • The 5th floor passed clearance for mold testing, except for rooms 550, 531, 556 and 557.

Dental School :

  • Sherlock has most of the new walls for the temporary bookstore framed and sheet rocked. 
    • They have installed the duct board and flex duct for the supply air and return air to the temporary bookstore area as well as the existing adjacent rooms.
  • Trison Constructors completed the demolition of the existing oven duct work and began the installation of the new ducts in the basement and on the 1st Floor. 
  • Needham Roofing removed approximately one half of the penthouse roof on the Clinic Building and removed most of the debris from the roof.
  • Securitas ran coax cables for the intercom system to the front desk located in the Clinic Building , 1st Floor lobby.

Landscaping Project:

  • Two additional irrigation stations have been completed on the Perdido St. side of the MEB yesterday.
    • Weather prevented crews from working today.